OpenStudy (whimsical):

Please help Chris takes a bus to school every day in a 5-day week. The bus journey consists of two intermediate stops. The probability of delay at stop A is 4/5 while the probability of delay at stop B is 2/5. The delays are independent of each other. If the bus is delayed at either stop, Chris will be late for school. Given that Chris was late on Monday, find the probability he was late exactly three times in a week. i know its conditional probability, however i am not sure how to get the probability of Chris being late on Monday.

OpenStudy (anonymous):

given that chris is late on monday means you have to compute the probability that he is late 2 out of the remaining 4 days

OpenStudy (anonymous):

probability he is delayed is pretty damned good considering it is already \(\frac{4}{5}\)at one stop probability he is delayed at one or the other or both stops is \(\frac{4}{5}+\frac{2}{5}-\frac{4}{5}\times \frac{2}{5}=\frac{22}{25}\)

OpenStudy (anonymous):

two out of the four days is computed by \[\dbinom{4}{2}(\frac{22}{25})^2\times (\frac{3}{25})^2\]

OpenStudy (whimsical):

ok thank you i think i understood